Program of the 38-th Symposium on Ring Theory and Representation Theory(2005)

September 2 (Friday)

[09:45 – 10:30] Ohnuki, Yosuke (Tokyo Univ. of Agr. and Tech.) Stable equivalences related with syzygy functors

[10:45 – 11:30] Nagase, Hiroshi (Nara National College of Tech.) Hochschild cohomology of stratified algebras

[11:45 – 12:30] Nishida, Kenji (Shinshu Univ.) Linkage and duality of modules

[13:45 – 14:45] Ariki, Susumu (Kyoto Univ.) Algebras arising from Schur-Weyl type dualities (I)

[15:00 – 15:45] Yanagawa, Kohji (Osaka Univ.) Castelnuovo-Mumford regularity for complexes and weakly Koszul modules

[16:00 – 16:45] Motose, Kaoru (Hirosaki Univ.) Integral group algebras and cyclotomic polynomials

[17:00 – 18:00] Rouquier, Raphaël (Université Denis Diderot - Paris 7) Broué’s conjecture: methods and results (I)

September 3 (Saturday)

[09:45 – 10:30] Iyama, Osamu (Univ. of Hyogo) Orthogonality of subcategories

[10:45 – 11:30] Takahashi, Ryo (Meiji Univ.) A generalization of n-torsionfree modules

[11:45 – 12:30] Koshitani, Shigeo (Chiba Univ.) Broué’s abelian defect group conjecture

[13:45 – 14:45] Ariki, Susumu (Kyoto Univ.) Algebras arising from Schur-Weyl type dualities (II)

[15:00 – 16:00] Rouquier, Raphaël (Université Denis Diderot - Paris 7) Broué’s conjecture: methods and results (II)

[16:15 – 17:00] Tachikawa, Hiroyuki (Univ. of Tsukuba) QF rings and QF associated graded rings

September 4 (Sunday)

[09:45 – 10:45] Rouquier, Raphaël (Université Denis Diderot - Paris 7) Broué’s conjecture: methods and results (III)

[11:00 – 12:00] Ariki, Susumu (Kyoto Univ.) Algebras arising from Schur-Weyl type dualities (III)